### v1

# CATArena: Evaluation of LLM Agents Through Iterative Tournament Competitions

Abstract
Large Language Model (LLM) agents have evolved from basic text generation to autonomously completing complex tasks through interaction with external tools. However, current benchmarks mainly assess end-to-end performance in fixed scenarios, restricting evaluation to specific skills and suffering from score saturation and growing dependence on expert annotation as agent capabilities improve. In this work, we emphasize the importance of learning ability, including both self-improvement and peer-learning, as a core driver for agent evolution toward human-level intelligence. We propose an iterative, competitive peer-learning framework, which allows agents to refine and optimize their strategies through repeated interactions and feedback, thereby systematically evaluating their learning capabilities. To address the score saturation issue in current benchmarks, we introduce CATArena, a tournament-style evaluation platform featuring four diverse board and card games with open-ended scoring. By providing tasks without explicit upper score limits, CATArena enables continuous and dynamic evaluation of rapidly advancing agent capabilities. Experimental results and analyses involving both minimal and commercial code agents demonstrate that CATArena provides reliable, stable, and scalable benchmarking for core agent abilities, particularly learning ability and strategy coding.
[…]
1 Introduction
With the rapid evolution of agents powered by large language models (LLMs), their capabilities have far surpassed simple text generation. By actively invoking external tools, LLM agents have significantly expanded the boundaries of artificial general intelligence (AGI). These agents are now able to autonomously complete complex, multi-step tasks that are previously considered beyond their reach, such as developing software project (Manish, 2024; Hu et al., 2025b), intelligently performing strategic planning (Belle et al., 2025), and learning user preference (Gao et al., 2024).
Existing benchmarks mainly focus on end-to-end performance in specific tasks, such as code generation (Yang et al., 2024), AI research (Nathani et al., 2025), and GUI automation (Wang et al., 2024). These benchmarks provide detailed observations and analyses of LLM agents’ abilities within particular scenarios and have driven significant progress in the field. However, there are important limitations to these approaches. First, the scores obtained in these end-to-end benchmarks only reflect performance on specific tasks, whereas an agent’s overall capability is composed of multiple fundamental skills working together. Second, the absolute scores in these benchmarks, which are typically based on objective correctness, have an upper bound. As agents become increasingly powerful, maintaining and updating these benchmarks requires additional expert-level annotation, and the level of required expertise continues to rise. In light of these challenges, there is an urgent need for a quantifiable and continuously evolving benchmark that systematically measures and analyzes the fundamental sub-abilities of agents.
Previous research has shown that self-learning is an essential ability for agents to achieve human-level intelligence (Gao et al., 2025; Zhu et al., 2025). Beyond self-learning, agents, similar to humans, also engage in peer learning, which enables collective evolution through interactions and shared experiences (Liu et al., 2024). During this evolutionary process, agents receive feedback from their environment and continually improve themselves. This capacity for learning and adaptation is indispensable for LLM agents, as it prepares them for ongoing evolution and more complex challenges. To systematically evaluate this crucial ability, we propose an iterative peer-learning-based competitive framework for LLM agents. In each iteration, agents are required to revise and update their strategies based on the outcomes and policies observed in previous rounds of competition. After every update, the agent policy codes are executed and competed against each other, generating dynamic performance rankings. Through this peer-learning architecture, we gain valuable insights into the learning abilities of LLM agents.
Building on this peer-learning framework, we introduce CATArena (Code Agent Tournament Arena) [^1] Code of CATArena is available in https://github.com/AGI-Eval-Official/CATArena., which utilizes four open-ended, rankable games. These games, including both board games and card games, provide LLM agents with a peer-learning environment and unlimited upper bound for improvement. They enable agents to continually improve and compete, ensuring that the evaluation framework remains challenging as agent capabilities grow. Furthermore, our competitive arena is inherently extensible and can be readily adapted to other types of open-ended, rankable tasks, facilitating the assessment of core agent abilities in new domains. As agent capabilities continue to advance, CATArena can evolve by incorporating tasks with greater complexity and discrimination, thereby supporting ongoing evaluation without the need for expert-level human annotation.
In our experiments, we conduct comparative performance evaluations and data analysis conducted on our self-developed minimal code agent and state-of-the-art commercial code agents. CATArena consistently provides stable and reliable benchmarks for assessing both agent capabilities and the agentic potential of the underlying LLMs. Within the peer-learning framework, we design general scoring metrics to systematically assess the fundamental abilities of participating agents, including their learning ability. Our experiments demonstrate that the strategy coding tasks applied in CATArena are fundamentally different from traditional LLM reasoning tasks. This represents a novel evaluation dimension that has not been addressed in previous work. Additionally, we analyze characteristics of CATArena, demonstrating its reliability and extensibility as a benchmarking platform.
In summary, our contributions are as follows:
  • Iterative Peer-learning-based Competitive Framework: We propose a novel framework that leverages iterative peer-learning and competition to evaluate the learning abilities of LLM agents. Agents continuously revise their strategies based on feedback and outcomes from previous rounds, aligning agent evolution with human evolution.
  • CATArena Benchmark: We introduce CATArena, a tournament-style benchmark for evaluating the basic capabilities of LLM agents using a diverse set of open-ended games, including board and card games. CATArena provides an unlimited upper bound for agent improvement and supports extensible evaluation across diverse, open-ended tasks.
  • Comprehensive Agent Evaluation: We design general and systematic evaluation matrices and conduct comparative experiments and analyses between our minimal code agent and state-of-the-art commercial agents, demonstrating the reliability, stability, and extensibility of the CATArena.
2 Related Work
Learning Ability.
Learning ability is crucial for LLM agents, as it enables continual adaptation and improvement in dynamic and complex environments. Recent studies have shown that self-learning methods, such as self-refinement (Madaan et al., 2023; Shinn et al., 2023), allow models to enhance their outputs through iterative feedback, while environmental feedback further supports continual learning (You et al., 2024). In addition to self-learning, peer-learning has also been increasingly recognized, with approaches encouraging agents to learn from others’ reasoning processes and shared experiences (Liang et al., 2024; Luo et al., 2025). These diverse learning mechanisms have led to notable advances in tasks such as code generation, complex reasoning, and collaborative problem-solving. In the context of LLM-driven agents, learning ability represents a critical capability that supports effective adaptation and enables agents to tackle increasingly complex tasks and evolving challenges (Zhu et al., 2025; Gao et al., 2025). Despite the progress, systematic evaluation of how agents learn from each other remains underexplored, highlighting the need for benchmarks that capture both self-learning and peer-learning abilities.
Evaluation on Agents.
Recent benchmarks primarily assess LLM-driven agents on end-to-end, task-specific abilities. Code-based evaluations such as GitTaskBench (Ni et al., 2025), SUPER (Bogin et al., 2024), ProjectEval (Liu et al., 2025), SWE-PolyBench (Rashid et al., 2025), RedCode (Guo et al., 2024), SWT-Bench (Mündler et al., 2025), InfiAgent-DABench (Hu et al., 2024), and DA-Code (Huang et al., 2024) focus on large-scale software development, code security, bug fixing, and data science tasks. Other works extend agent evaluation to research (Du et al., 2025), real-world tool use (Yao et al., 2024), and assistant scenarios (Mialon et al., 2024). While some benchmarks explore agent-vs-agent evaluation (Zhuge et al., 2024), most rely heavily on human annotation and objective correctness, leading to upper bounds and saturation as agent capabilities advance.
Open-ended Tasks.
To address these limitations, recent work has leveraged open-ended, rankable tasks. Benchmarks such as GameBench (Costarelli et al., 2024), lmgame-Bench (Hu et al., 2025a), GAMEBot (Lin et al., 2024), and card game evaluations (Wang et al., 2025) assess LLMs’ strategic reasoning through diverse games. Frameworks like Game Reasoning Arena (Cipolina-Kun et al., 2025), GVGAI-LLM (Li et al., 2025), ZeroSumEval (Alyahya et al., 2025), TextArena (Guertler et al., 2025), and MCU (Zheng et al., 2025) further extend evaluation to multi-turn reasoning, spatial adaptability, natural language interaction, and open-ended tasks. However, these benchmarks mainly focus on reasoning skills and do not systematically evaluate agents’ learning abilities or coding strategies. Our analysis shows that measuring learning ability and coding strategy is fundamentally distinct, and both are essential for advancing agent intelligence. Additionally, in human board game competitions, variant rules such as Chess960 (FIDE, 2023) and Six-plus Hold’em (sixplusholdem, 2025) are often introduced to reduce memorization and encourage creativity. Notably, these variant rules have received relatively less attention in model evaluation.
3 CATArena
Figure 1: Overview of the evaluation framework and CATArena. The evaluation framework adopts an iterative peer-learning based competitive process. In round 1, LLM agents develop initial strategies via coding. These strategies are matched in a tournament arena, producing rankings and logs. In each subsequent round, agents analyze previous codes and logs, refine their strategies, and compete again. CATArena includes four open-ended and rankable games to cover diverse settings. Based on the tournament results from all rounds, a carefully designed scoring matrix and evaluation scheme are used to robustly quantify various agent abilities.
3.1 Iterative Peer-Learning based Competitive Framework
As shown in Figure 1, we propose an iterative peer-learning-based competitive framework, where CATArena evaluates code agents through a two-phase workflow: initial strategy development and iterative improvement. The initial phase assesses each agent’s ability to independently implement a baseline strategy based on the game code, while the iterative phase focuses on the agent’s learning ability.
Initial Development (Round 1). In this stage, each agent receives the game code and a sample AI implementation. Without external guidance, each agent must develop its own strategy to participate in the tournament. This phase primarily examines the agent’s strategy coding ability and establishes a baseline for subsequent evaluation.
Iterative Improvement (Rounds $n>1$). After the first round, all strategies submitted are evaluated through a tournament (round-robin format for symmetric games, batch-based competition for asymmetric games). Comprehensive competition logs are generated, recording rankings, win counts, and move histories for all matches. In subsequent rounds, agents are provided with the game code, previous round submissions from all participants, and these detailed logs. Agents must analyze these resources from previous rounds to adapt and improve their own strategies. This phase assesses the agent’s learning ability through repeated cycles of analysis and refinement.
This iterative evaluation framework of CATArena enables a granular assessment of both basic coding skills and advanced learning capabilities, supporting a robust and scalable measurement of code agent performance.
Table 1: Overview of game arenas and representative variants in CATArena.
Game Symmetry Type Players Variant
Gomoku Board 22 Forbidden points; dual three-in-a-row
Texas Hold’em Card $\geq 8$ Card removal; swapped hand ranks
Chess Board 22 Chess960; forbidden/special moves
Bridge Card 44 Card exchange
    For Bridge, symmetry is defined by assigning identical agent strategies to both teammates.
3.2 Games and Variants
Building on the tournament-based evaluation framework, CATArena deploys four distinct game arenas, each selected to test the strategic reasoning and coding capabilities of code agents across varying levels of complexity and interaction patterns. These arenas include competitive and cooperative settings, as well as symmetric and asymmetric game structures, thus enabling a diverse analysis of agents’ strategy coding abilities and learning patterns.
In addition to standard rules, each game is extended with thoughtfully designed variants that introduce novel or altered mechanics, inspired by real-world adaptations such as Fischer Random Chess (Chess960) (FIDE, 2023). Like human competition, the variant rules encourage strategy generalization and penalize rote memorization, as most models are trained on card and board game data. Table 1 provides an overview of the selected games and their respective variants.
3.3 Tournament Format and Scoring System
After the completion of all $N$ development rounds, CATArena conducts a comprehensive tournament to quantitatively evaluate agent strategies and compute performance metrics. A total of $T$ agent models participate, each contributing strategies in every round of development. Tournament formats are tailored to game types: for symmetric games, all strategies engage in a round-robin cycle, ensuring exhaustive pairwise competition; for asymmetric games such as Texas Hold’em, strategies are grouped into batches and compete in multi-agent matches. To mitigate randomness, all matches are repeated multiple times, and results are averaged for robust evaluation.
Scores are recorded in a scoring matrix $W\in\mathbb{R}^{(TN)\times(TN)}$, where $W_{i,j}^{n,m}\in[0,1]$ denotes the score obtained by agent $i$’s strategy in round $n$ against agent $j$’s strategy in round $m$. When $n=m$, the notation simplifies to $W_{i,j}^{n}$; similarly, when $i=j$, it is denoted as $W_{i}^{m,n}$. This scoring system enables fine-grained, quantitative analysis of agent performance in both individual and iterative development stages. For asymmetric games, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches. The tournament format is provided in the Appendix A.
3.4 Evaluation Metrics
Based on the scoring matrix $W$, we design a set of evaluation metrics to quantitatively assess the key capabilities of code agents. Specifically, our metrics are constructed to measure three core capabilities: strategy coding, learning, and generalizability. In the following sections, we define these metrics using symmetric games as examples. For asymmetric games, the evaluation principles remain consistent. The calculation of the scoring matrix $W$ is provided in Appendix B.
Strategy Coding.
Strategy coding measures the agent’s fundamental ability to abstract game strategies into reproducible algorithms and implement them as executable code, which is fundamentally different from general reasoning and strategic planning abilities. In CATArena, this metric evaluates how effectively an agent can independently develop a baseline strategy for the game environment and compete against other agents in the initial development stage.
For each agent $i$, strategy coding is quantified by the average score obtained against all other agents in the first round:
$S_{i}=\mathrm{avg}_{j\neq i}(W_{i,j}^{1}).$
This metric serves as the foundational benchmark for code agent evaluation in CATArena.
Learning Ability.
The learning capability of a code agent captures its ability to leverage historical information and opponent behaviors to improve its own performance.
Global Learning assesses an agent’s overall improvement in strategy quality. This metric evaluates the relative performance of agent $i$’s strategies against all strategies from all agents and rounds, and measures the average progress made compared to its initial baseline. It serves as the primary indicator of learning ability.
Formally, for agent $i$, global learning is defined as:
$L_{i}=\mathrm{average}_{n=2}^{N}\left(G_{i}^{n}-G_{i}^{1}\right),$
where $G_{i}^{n}$ represents the global performance of agent $i$’s strategy from round $n$:
$G_{i}^{n}=\mathrm{average}_{(i,n)\not=(j,m)}\left(W_{i,j}^{n,m}\right).$
This metric captures the agent’s ability to learn and adapt over multiple rounds, reflecting its progress in a comprehensive competitive landscape.
Counter-Adaptation measures an agent’s targeted learning ability, reflecting its capacity to achieve improved results against opponents in successive rounds. For agent $i$, the counter-adaptation score is defined as the average improvement in scores against other agents from round n1n-1 to round $n$ $(n\geq 2)$:
$C_{i}=\mathrm{average}_{n=2}^{N}\left(A_{i}^{n}-B_{i}^{n-1}\right),$
where the advance score $A_{i}^{n}$ and base score $B_{i}^{n-1}$ are defined as:
$A_{i}^{n}=\mathrm{average}_{j\neq i}\left(W_{i,j}^{n,n-1}\right),B_{i}^{n-1}=\mathrm{average}_{j\neq i}\left(W_{i,j}^{n-1}\right).$
Here, $A_{i}^{n}$ represents agent $i$’s average performance in round $n$ against the strategies submitted by other agents in the previous in round (n1n-1). The base score $B_{i}^{n-1}$ denotes agent $i$’s average performance in round n1n-1 against those same opponents. This comparison isolates the agent’s targeted adaptation from one round to the next.
Self-improvement evaluates an agent’s capacity to genuinely enhance its strategies over successive rounds of development. This metric reflects whether newly developed strategies can consistently outperform the agent’s own previous versions.
We quantify self-improvement by calculating the Pearson correlation Pearson (1896) between the round index and the agent’s average scores across rounds. For agent $i$, the self-improvement score is defined as
$\mathrm{SI}_{i}=\mathrm{Pearson}\left([1,\cdots,N],\ [S_{i}^{1},\cdots,S_{i}^{N}]\right).$
Here, $S_{i}^{n}$ denotes the average score of agent $i$’s strategy from round $n$ against its own strategies from other rounds
$S_{i}^{n}=\mathrm{average}_{m\neq n}\left(W_{i}^{n,m}\right).$
A higher self-improvement score indicates a stronger ability to iteratively refine and upgrade strategies throughout the development process.
Generalizability.
Generalizability measures an agent’s ability to comprehend and adapt to novel or altered game rules that differ from those encountered during training or prior experience. This metric specifically evaluates the agent’s capacity to generalize beyond previously seen environments, focusing on handling new or modified scenarios. For agent $i$, the generalizability score is defined as:
$U_{i}=B_{i}^{1;\mathrm{Variants}}-B_{i}^{1;\mathrm{Standard}},$
where $B_{i}^{1;\mathrm{Variants}}$ and $B_{i}^{1;\mathrm{Standard}}$ denote the base scores of agent $i$ in the first round under variant and standard rule settings, respectively. A higher value of $U_{i}$ indicates stronger generalizability, reflecting the agent’s ability to effectively develop and apply strategies for previously unseen tasks.
4 Experiments
4.1 Experimental Setups
Participants.
In our experiment, we employ three types of agents: (1) Minimal Agents (LLM + ADK Framework) [^2] https://github.com/AGI-Eval-Official/Minimal-CodeAgent: A baseline agent developed with the Agent Development Kit (ADK) Python toolkit. We provide essential tools, including file manipulation, bash scripting, and Python execution, for the ADK code agent. On this foundation, we integrate state-of-the-art LLMs to systematically compare their core competencies as code agents for strategy implementation. (2) Commercial Code Agents: State-of-the-art and commercial CLI-based agents (e.g., Claude Code, CodeX, Gemini-CLI, Qwen-Coder) are included for benchmarking. These agents feature advanced integration with various command-line interfaces, tools, and LLMs, resulting in enhanced overall capabilities. These agents serve as leading solutions in code agent development and provide valuable reference points for future research. (3) LLM-Player: In this control setting, LLMs directly output game moves without generating code. For each turn, the LLM receives the game rules, current state, and history, and returns the next action. This approach is specifically designed to assess the inherent strategic and reasoning capabilities of LLMs. Detailed agent parameter settings and model selections are presented in Appendix C.
Table 2: Agent Specifications and Open-Source Status.
Agent Type Agent Framework Model Agent OSS LLM OSS
Minimal basic code tools with ADK framework DeepSeek-3.1 (DeepSeek-AI, 2024)
Qwen3-Coder-480B (Team, 2025c)
Doubao-Seed-1.6 (Team, 2025d)
GPT-5 (OpenAI, 2025b)
Claude-4-Sonnet (Anthropic, 2025a)
Gemini-2.5-pro (Team, 2025a)
Commercial Gemini-CLI (Google, 2025) Gemini-2.5-pro (Team, 2025a)
Claude-Code (Anthropic, 2025b) Claude-4/3.7 Hybrid (Anthropic, 2025a)
CodeX (OpenAI, 2025a) GPT-5 (OpenAI, 2025b)
Qwen-Coder (Team, 2025b) Qwen3-Coder-480B (Team, 2025c)
Other LLM-Player Agents’s Corresponding LLM N/A N/A
Tournaments.
All experiments are conducted under two main tournament settings: (1) a comparison among minimal agents equipped with different LLMs ($T_{1}=6$), and (2) a comparison between the best-performing minimal agent and a set of commercial code agents ($T_{2}=5$). To reduce the impact of randomness on strategy generation, each tournament is repeated for four times, and all reported metrics are averaged over the four runs. Each tournament consists of $N=4$ rounds of iterative development. To further mitigate stochastic effects in competition outcomes, every entry in the scoring matrix $W$ is estimated by repeated matches. Detailed tournament prompts are listed in Appendix K.
We report the detailed scoring policy, generation configs, and repetition experiments in Appendix D. It is noteworthy that agents tend to generate different codes in repeated experiments, but their rankings are relatively stable.
4.2 Main Results
(a) Trends of global performance scores $G_{i}^{n}$.
(b) Quantitative comparison of learning abilities.
Figure 2: Visualization of agents’ learning patterns and scores.. For clarity, we use family names to represent LLM models instead of their full names. The results for other games are in Appendix E.
Table 3: Main LeaderBoard of CATArena. We conduct two groups of tournaments between minimal agents (rank from 1 to 6) and commercial agents (rank from 1 to 5), and report the average ranking across all four tasks. Metrics include S.C. (Strategy Coding), G.L. (Global Learning), and G.A. (Generalizability).
Agents Standard Variant G.A. $\downarrow$
S.C. $\downarrow$ G.L. $\downarrow$ S.C. $\downarrow$ G.L. $\downarrow$
Minimal
Claude-4-Sonnet 1.25 2.5 1.75 2.75 5.00
DeepSeek-Chat 5.75 2.75 4.25 2.75 2.75
Doubao-Seed 3.75 4.75 3.75 4.50 2.75
Gemini-2.5-Pro 3.25 3.75 3.25 2.75 3.25
GPT-5 3.75 3.50 3.00 3.75 2.25
Qwen3-Coder 2.25 3.75 3.00 4.5 4.75
Commercial
best ADK 3.25 2.25 2.00 3.75 2.50
Claude-Code 2.50 3.75 2.50 2.75 3.25
CodeX 2.25 2.75 3.00 3.00 3.25
Gemini-CLI 3.50 2.25 3.00 4.00 2.00
Qwen-Coder 3.00 3.75 4.00 1.25 3.25
Learning Ability.
Figure 2(a) visualizes the global performance scores $G_{i}^{n}$ revealing overall trends in agent strategies across iterations. Some agents, such as minimal agents driven by claude, exhibit a clear upward trajectory over multiple rounds, demonstrating strong learning capability. However, the performance of most agents remains unstable, and no obvious trend is observed. To explicitly illustrate the learning ability of each agent, we design a quantitative scoring method for global learning and introduce two additional learning modes: counter-adaption and self-improvement. The quantitative results for these three abilities are presented in Figure 2(b). Compared to Figure 2(a), these scores offer a more intuitive comparison of the agents’ learning strengths. This decomposition provides deeper insights into the mechanisms underlying agent learning. When both counter-adaption and self-improvement scores are positive, it indicates that the agent can effectively learn from both its opponents and itself, resulting in a positive global learning score. Compared to minimal agents, commercial agents exhibit stronger learning capabilities. Learning ability results for other games and a case study of strategy iteration mechanisms are provided in Appendix E. Through an analysis of the consistency of agent actions in endgame states, we observe that a majority of agents indeed learn from the code generated in the previous round, leading to increasing consistency. This phenomenon is most prominent in the Hold’em environment, possibly because its strategies are relatively simple and easier to learn.
Ability Evaluation.
We conduct two sets of tournaments: (1) minimal agents equipped with different LLMs, (2) the best-performing minimal agent against commercial code agents. For each tournament, we report the average ranking and scores for three core agent capabilities. The main leaderboard and main results of CATArena are summarized in Table 3 and Table 4. These results reveal the following key observations:
Observation 1: The performance gap among LLMs is more pronounced in minimal agents compared to commercial agents. Table 3 shows that Claude-4-Sonnet achieves the highest score among minimal agents, while the rankings of other LLMs are more dispersed. In contrast, commercial agents driven by the same LLMs exhibit much closer average rankings, with all agents scoring around 2.5 out of 5, indicating a reduced performance gap. Moreover, commercial agents demonstrate performance levels similar to the best-performing minimal agent. This suggests that the underlying agent framework can significantly influence how effectively an LLM’s capabilities are utilized, as commercial agents are often optimized for specific models.
Observation 2: The participating agents display different ranking orders across various capabilities. The tournament results reveal that the relative rankings of agents change depending on the specific core ability being tested. The ranking of these abilities provides a decomposition of end-to-end performance, offering insights for further optimization of both LLMs and agent frameworks.
Observation 3: Agents exhibit varied performance distributions across different tasks. The results indicate that agents’ performance is not uniformly distributed across all tasks, which is mainly attributed to the distinct nature and difficulty of the four tasks. In the variant tasks, the performance gap among agents is more pronounced, likely because the game rules and strategies are less familiar to agents.
[…]
Games Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant
[…]
Minimal
Claude-4-Sonnet 0.88 -0.447 0.78 -0.156 -0.14 0.58 0.118 0.13 0.110 -0.45 0.79 0.174 1.0 0.047 0.005 0.90 -0.170 0.65 0.018 -0.55
Deepseek-Chat 0.23 0.027 0.38 0.077 0.13 0.01 0.010 0.00 -0.022 -0.01 0 0 0.10 0 0.10 0 0 0.10 0.049 0.10
Doubao-Seed 0.33 -0.192 0.72 -0.302 0.46 0.04 -0.035 0 0 -0.04 0.2 -0.033 0.45 -0.516 0.40 0.58 -0.337 0.10 0.034 -0.46
Gemini-2.5-Pro 0.25 -0.066 0.00 0.173 -0.12 0.01 0.020 0.00 0.078 -0.01 0.90 -0.195 0.60 -0.049 -0.30 0.58 -0.147 0.90 0.003 0.46
GPT-5 0.48 0.062 0.76 -0.019 0.18 0.16 0.102 0.87 -0.050 0.71 0.47 -0.095 0.10 0.293 -0.02 0.38 -0.525 0.45 -0 0.24
Qwen3-Coder 0.85 -0.523 0.36 -0.089 -0.50 0.20 0.038 0.00 0.003 -0.20 0.65 0.032 0.76 -0.230 -0.19 0.58 -0.187 0.80 -0.532 0.21
Commercial
best ADK 0 0.075 0.75 -0.022 0.88 0.07 0.073 0.46 0.030 0.39 0.25 0.295 0 0.361 -0.25 0.91 -0.110 1.00 -0.342 -0.47
Claude-Code 0.78 -0.322 0.66 0.194 -0.28 0.01 0.100 0 0.105 -0.001 1.00 -0.240 0.93 -0.139 -0.04 0.56 -0.158 0.44 -0.226 0.03
CodeX 0.47 0.454 0.69 -0.095 0.34 0.72 0.050 0.17 0.067 -0.55 0.75 0.098 0.50 0.285 -0.25 0.38 0.033 0.34 0.064 0.09
Gemini-CLI 0.31 0.260 0.19 0.172 0.0 0.13 0.050 0.37 -0.007 0.24 0.01 0.254 0.83 -0.286 0.79 0.38 0.395 0.38 -0.154 0.16
Qwen-Coder 0.94 -0.054 0.22 -0.530 -0.94 0.07 0.058 0 0.105 -0.007 0.49 0.157 0.25 0.556 -0.25 0.28 -0.204 0.34 0.039 0.19
4.3 Effectiveness of CATArena
We design a series of experiments to demonstrate the effectiveness of CATArena.
Comparison between Agents and LLM-Players.
The primary task in CATArena is strategy coding, which relies on the underlying coding capabilities of LLMs. We posit that reasoning over code to develop strategies is fundamentally different from direct reasoning during gameplay. To validate this distinction, we compare agent-developed strategies with the LLM-Player baseline (see Appendix F). Our results show that current agents primarily implement simple rule-based algorithms, indicating substantial room for advancement in agents’ strategy coding abilities. CATArena fully leverages this non-saturation, enabling sustainable iterative peer-learning.
To further analyze the similarities and differences between agent-implemented code strategies and those of LLM-Players, we ask agents’ code and LLM-Player to select the next action on endgame states. Figure 3 illustrates the action consistency between agents’ code and LLM-Players in Chess. Surprisingly, the strategies encoded in agent code differ significantly from those inferred directly by the LLM, even if they are from the same model. Meanwhile, strategies produced by different agents and different LLMs also show notable similarities. This indicates that strategy coding and reasoning in LLMs are distinct capabilities. We report results of other tasks are demonstrated in Appendix G. CATArena evaluates the strategy coding ability of agents rather than their reasoning ability, thereby filling a gap in previous benchmarks. The relationship between strategy coding ability and LLM-based strategy reasoning remains unclear and requires further investigation.
Figure 3: Action consistency between agents’ code and LLM-Players on Chess endgames.
Collective Learning Trends Among Agents.
Table 5: Collective learning trends of agents across different tasks in CATArena. $\text{DIS}_{\text{range}}$ and $\text{DIS}_{\text{std}}$ represent the Pearson correlation coefficients of the range and standard deviation of agent performance scores over four rounds, reflecting the similarity and dispersion of agent strategies. $\text{Trend}_{\text{mean}}$ denotes the Pearson correlation between the mean agent performance and the round number, indicating the overall trend of group improvement.
Gomoku Hold’em Bridge Chess
StdV. VarV. StdV. VarV. StdV. VarV. StdV. VarV.
$\text{DIS}_{\text{std}}$ -0.05 0.15 -0.81 -0.80 -0.82 -0.57 0.55 -0.04
$\text{DIS}_{\text{range}}$ -0.16 0.44 -0.80 -0.76 -0.54 -0.33 -0.08 0.16
$\text{Trend}_{\text{mean}}$ 0.42 -0.02 0.75 0.67 0.24 -0.10 -0.74 -0.79
We analyze the collective learning dynamics of agents across tasks, as presented in Table 5. The metrics $\text{DIS}_{\text{range}}$ and $\text{DIS}_{\text{std}}$ report the Pearson correlation between the standard deviation and range of agent performance scores over four rounds ($B_{i}^{n},n=1,2,3,4$). The higher similarity in performance scores (i.e., $\text{DIS}>0$) suggests that agents can learn effective strategies more readily, indicating lower task difficulty. Based on these results, the relative difficulty ranking of tasks in CATArena is Chess >> Gomoku >> Bridge >> Hold’em.
[…]
Additional Results.
CATArena’s iterative peer-learning framework is easily extensible to new tasks for evaluating other fundamental agent abilities. We demonstrate this by introducing a Machine Learning (ML) track and multi-lingual track, with experimental results provided in Appendix H and Appendix I, respectively. Experimental results indicate that current agents still exhibit substantial potential for improvement. As agents continue to advance, the open-ended task design and peer-learning evaluation framework of CATArena ensure that systematic assessment can be sustained over time. We also report agent cost in terms of token usage, time consumption, and generated code statistics for each agent, in Appendix J. Notably, Claude-4-Sonnet utilizes the most tools and tokens, and also develops a significantly larger amount of code. In contrast, GPT-5 achieves the best balance between token usage and performance. Efficient utilization of tokens and external tools remains an important research direction to advance the capabilities of LLM agents.
5 Conclusion
In this work, we address two fundamental challenges in LLM agent evaluation: the need for systematic measurement of learning ability, and the tendency of traditional benchmarks to become saturated as agent capabilities improve. To this end, we propose an iterative peer-learning-based competitive framework, enabling agents to continually revise and enhance their strategies through dynamic interaction and feedback. Building on this, we introduce CATArena, a tournament-style benchmark featuring open-ended and rankable board and card games. CATArena provides an environment with unlimited potential for agent improvement potential and extensible evaluation across new domains. Experimental results demonstrate that our framework reliably assesses core agent abilities, particularly learning ability and strategy coding, while ensuring stability and scalability. The open and flexible architecture of CATArena supports ongoing research and benchmarking for future intelligent agents.
Limitations.
The current evaluation in CATArena is limited to four games, which primarily assess agents’ learning ability and strategy coding. These scenarios do not encompass the full spectrum of potential LLM agent capabilities. In future work, we plan to introduce a wider variety of more complex tasks to evaluate agents’ learning and other abilities from different perspectives.
Appendix A Tournament Format and Scoring System
Table 6: Configs of tournament on four games.
Environment Code Agents LLM-player
Gomoku
Board size: $\times15$
Number of pairwise matches: 4 $\times$ 2
Swap black and white pieces after each match
Maximum time per move: 10 s
Board size: $\times15$
Number of pairwise matches: 2 $\times$ 2
Swap black and white pieces after each match
Maximum time per move: 600 s
Texas Hold’em
Max players: 12
Rounds: 100
Random shuffle seat after each round
Initial chips: 2000
Blind increase every 24 hands
Max hands per round: 720 or until winner decided
Maximum time per move: 3 s
Max players: 12
Rounds: 10
Random shuffle seat after each round
Initial chips: 2000
Blind increase every 24 hands
Max hands per round: 720 or until winner decided
Maximum time per move: 1000 s
Bridge
Number of pairwise matches: 12 $\times$ 2
Swap directions of open/closed rooms
Use same deck for each pair of match
Maximum time per move: 10 s
Number of pairwise matches: 12 $\times$ 2
Swap directions of open/closed rooms
Use same deck for each pair of match
Maximum time per move: 200 s
Chess
Number of pairwise matches: 8 $\times$ 2
Swap black and white pieces after each match
Maximum moves per game: 200
Maximum time per move: 10 s
Number of pairwise matches: 2 $\times$ 2
Swap black and white pieces after each match
Maximum moves per game: 200
Maximum time per move: 600 s
We list the basic settings for each games in Table 6.
For each game, we ensure that the number of pairwise matches among code agents allows the final results to stabilize, i.e., for each game, the L1-norm fluctuation of the scoring matrix $W$ is less than 5%.
For LLM-players, since their reasoning time is relatively long for most games, we reduce the number of repeated experiments. Note that our comparison with LLM-players is only a qualitative analysis of the differences between LLM-player and corresponding coding agent. The exploration of LLM-players’ results is not the focus of this paper; refer to the main text for details.
Appendix B Evaluation Metric Calculation
Table 7: Scoring rules of tournament on four games.
Environment Scoring Metric
Gomoku Pairwise match scoring: Win = 1 point,  Draw = 0.5 point,  Lose = 0 point.
Texas Hold’em Multi-agent batches: score is the average win rate across all tournaments participated.
Bridge 20 VP system: Two opposing pairs’ scores sum to 20, Final score divided by 20, ensuring each pair’s score $\in[0,1]$.
Chess Pairwise match scoring: Win = 1 point,  Draw = 0.5 point,  Lose = 0 point.
To evaluate the basic capabilities of LLM agents, we define metrics for each applied game. Let $N$ be the number of rounds and $K$ be the number of participating agents. We construct a matchup matrix
$W\in\mathbb{R}^{(N\cdot K)\times(N\cdot K)},$
where the generic element $W_{i,j}^{n,m}$ denotes the score when agent $i$ from round $n$ plays against agent $j$ from round $m$. We abbreviate $W^{n}_{i,j}$ for same-round comparisons ($n=m$) and $W_{i}^{n,m}$ for self-comparisons across rounds ($i=j$). Diagonal entries $(n,i)=(m,j)$ are ignored.
For asymmetric games, pairwise results are not feasible; instead, batch-based tournaments are used and the score matrix records the win rates of multi-agent matches. For each batch, we obtain a single score group $W_{i_{1},i_{2},\ldots,i_{BS}}^{n_{1},n_{2},\ldots,n_{BS}}$, where $BS$ is the batch size. We conduct three types of experiments to accommodate different metric calculations: (1) $W_{i}^{1,2,\ldots,N}$, where the same agent’s strategies from $N$ rounds compete against each other, used to compute self-improvement metrics $S_{i}^{n}$; (2) $W_{i_{1},i_{2},\ldots,i_{T}}^{n}$, where all agents in the same round compete, used to calculate the base score $B$; (3) All $N\times T$ agent strategies are randomly shuffled and grouped for competition (with $BS=12$ in our experiments), used to compute both the global score $G_{i}^{n}$ and advanced score $A_{i}^{n}$.
Scoring rules for the four games are summarized in Table 7.
Appendix C Generation Configs
For all LLMs used in our work, we set temperature to be $0.1$, max token identical to their official APIs’ setting. We set top-p to $1.0$, Top-k to be $100$, and presence penalty to be default to the API.
Additionally, both Claude-4-Sonnet and DeepSeek-3.1 occasionally encounter tool call issues that result in no code being generated, as frequently reported by the community. If such errors occur three times in a row, we substitute Claude-4-Sonnet with Claude-3.7-Sonnet and DeepSeek-3.1 with DeepSeek v3.
For LLM-players, considering the uncertainty in model output formats, we allow up to three retries. The prompt of LLM-players are in arena’s code and not present in paper considering its excessive length.
Appendix D Repetition Experiments
Table 8: Standard deviation of ranking in Round 1 and Round 2 with repeating 4 times.
Games Average Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Round 1
Minimal
Claude-4-Sonnet 0.80 0.91 1.58 0.43 1.12 0.71 0.50 1.64 0.00 0.87
Deepseek-Chat 0.72 0.81 0.83 0.87 0.83 1.22 1.22 0.71 0.00 0.43
Doubao-Seed 1.58 0.90 1.87 1.87 1.09 0.43 2.06 0.87 1.30 0.43
Gemini-2.5-Pro 1.24 1.23 1.50 1.30 1.12 0.83 1.12 1.66 1.22 1.12
GPT-5 0.75 1.18 0.71 1.50 0.71 1.30 1.09 1.50 0.50 0.43
Qwen3-Coder 1.16 0.84 1.48 1.50 1.22 0.71 1.09 0.71 0.83 0.43
Commercial
Claude-Code 1.27 1.01 1.12 1.66 0.43 0.00 1.79 1.09 1.73 1.30
CodeX 0.76 0.57 0.83 0.50 0.43 0.50 1.09 1.30 0.71 0.00
Gemini-CLI 1.14 0.93 0.83 0.43 1.12 1.00 1.30 1.79 1.30 0.50
Qwen-Coder 1.01 0.95 1.22 1.48 1.12 0.00 0.87 0.83 0.83 1.50
Round 2
Minimal
Claude-4-Sonnet 0.81 0.55 1.48 0.50 0.83 0.43 0.50 0.83 0.43 0.43
Deepseek-Chat 0.94 1.03 1.30 1.09 0.87 0.83 1.09 1.48 0.50 0.71
Doubao-Seed 0.79 0.88 1.09 1.12 0.87 0.71 0.71 0.87 0.50 0.83
Gemini-2.5-Pro 1.30 1.28 1.66 1.92 1.00 1.22 2.06 1.12 0.50 0.87
GPT-5 0.89 1.22 1.00 1.58 0.43 1.66 1.64 1.22 0.50 0.43
Qwen3-Coder 0.82 1.02 1.58 1.30 0.83 0.43 0.43 1.50 0.43 0.83
Commercial
Claude-Code 1.20 1.04 1.09 1.41 1.22 0.43 1.64 1.09 0.83 1.22
CodeX 0.71 0.87 0.43 0.87 0.50 1.09 0.83 1.09 1.09 0.43
Gemini-CLI 0.88 1.10 0.83 1.64 0.83 0.83 1.41 1.48 0.43 0.43
Qwen-Coder 1.13 0.90 1.66 0.83 0.43 0.83 1.00 1.12 1.41 0.83
We report the results of repetition experiments($N=4$) on first two tournament round in Table 8.
From table, we observe that 1. The rankings of most agents remain relatively stable ,with ranking standard deviation changes of less than one. However, a few agents, such as Gemini-2.5-Pro and Claude-Code, exhibit greater fluctuations; 2. The rankings for standard games are more stable than those for variant games; 3. The results of the open source model are more stable than those of the closed source model, and commercial agents are more stable than minimal agents; 4. Additionally, we observe that agents do not consistently generate runnable code repositories across multiple development attempts. Even commercial agents occasionally fail to produce successful builds, which suggests that current code agents still need to improve their development stability.
Appendix E Learning Ability
E.1 Global Learning Trend
(a) Gomoku.
(b) Hold’em
(c) Bridge.
(d) Chess.
Figure 4: Trend of global performance score $G_{i}^{n}$ in Gomoku, Hold’em, Bridge and Chess.
As shown in Figure 4, we present the trends of global performance scores $G_{i}^{n}$ across four games, revealing distinct performance patterns for different models. In many cases, agents experience a sharp decline in performance during an intermediate round, which we interpret as a learning failure. Typically, such failures are recovered in the following round.
E.2 Detailed Learning Score
We list the detailed score of global learning, counter-adpatation learning and self-improvement in table 91011 respectively.
Table 9: Global Learning with Group-wise Average Rankings.
Models Avg. Ranking Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Minimal
Claude-4-Sonnet 2.50 2.75 -0.447 -0.156 0.118 0.110 0.174 0.047 -0.170 0.018
Deepseek-Chat 2.75 2.75 0.027 0.077 0.010 -0.022 0.000 0.000 0.000 0.049
Doubao-Seed 4.75 4.50 -0.192 -0.302 -0.035 0.000 -0.033 -0.516 -0.337 0.034
Gemini-2.5-Pro 3.75 2.75 -0.066 0.173 0.020 0.078 -0.195 -0.049 -0.147 0.003
GPT-5 3.50 3.75 0.062 -0.019 0.102 -0.050 -0.095 0.293 -0.525 -0.000
Qwen3-Coder 3.75 4.50 -0.523 -0.089 0.038 0.003 0.032 -0.230 -0.187 -0.532
Commercial
best ADK 2.25 3.75 0.075 -0.022 0.073 0.030 0.295 0.361 -0.110 -0.342
Claude-Code 3.75 2.75 -0.322 0.194 0.100 0.105 -0.240 -0.139 -0.158 -0.226
CodeX 2.75 3.00 0.454 -0.095 0.050 0.067 0.098 0.285 0.033 0.064
Gemini-CLI 2.25 4.00 0.260 0.172 0.050 -0.007 0.254 -0.286 0.395 -0.154
Qwen-Coder 3.75 1.25 -0.054 0.536 0.058 0.105 0.157 0.556 -0.204 0.039
Table 10: Counter-adaptation Score with Group-wise Average Rankings.
Games Avg. Ranking Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Minimal
Claude-4-Sonnet 3.75 3.75 -0.096 -0.075 0.001 -0.023 0.005 0.128 -0.042 -0.075
Deepseek-Chat 2.75 2.00 0.008 0.088 0.004 0.021 0.000 0.000 0.000 0.100
Doubao-Seed 4.25 4.50 0.063 -0.196 -0.023 -0.061 -0.008 -0.133 -0.196 0.083
Gemini-2.5-Pro 2.75 2.75 0.354 0.192 0.014 0.097 -0.132 -0.238 -0.038 0.021
GPT-5 4.75 3.25 0.038 -0.012 -0.086 0.019 -0.098 0.052 -0.154 -0.025
Qwen3-Coder 2.75 4.75 -0.092 0.029 0.000 -0.080 0.037 -0.008 0.025 -0.167
Commercial
best ADK 2.50 3.00 0.260 0.000 0.018 0.039 0.127 0.238 -0.094 -0.104
Claude-Code 3.25 1.88 -0.042 0.083 0.091 0.194 0.060 0.081 -0.104 0.047
CodeX 3.50 3.50 0.104 -0.089 -0.034 0.033 0.023 -0.090 0.010 0.052
Gemini-CLI 2.75 4.25 0.120 -0.021 0.031 -0.083 -0.027 -0.169 0.188 0.010
Qwen-Coder 3.00 2.38 -0.130 0.193 0.041 0.077 0.092 0.081 -0.062 -0.021
Table 11: Self-improvement Score with Group-wise Average Rankings.
Games Avg. Ranking Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant Standard Variant
Minimal
Claude-4-Sonnet 2.75 2.50 -0.103 -0.894 0.949 0.517 0.858 0.766 -0.848 0.478
Deepseek-Chat 3.25 3.38 0.000 0.893 -0.949 -0.747 0.000 0.000 0.000 0.000
Doubao-Seed 3.75 4.38 0.141 -0.686 0.075 0.000 -0.202 -0.775 -0.894 0.000
Gemini-2.5-Pro 3.63 2.00 0.400 0.775 -0.758 0.894 -0.598 -0.240 -0.775 0.913
GPT-5 4.88 4.50 -0.897 -0.949 -0.205 -0.050 -0.767 0.485 -0.775 -0.390
Qwen3-Coder 2.75 4.25 -0.400 0.161 0.668 0.202 0.473 -0.400 -0.258 -0.775
Commercial
best ADK 2.25 3.50 0.400 0.956 0.835 -0.614 0.738 0.546 -0.207 -0.726
Claude-Code 4.00 2.25 -0.230 0.969 0.346 0.904 0.113 0.537 -0.730 -0.225
CodeX 2.00 3.25 0.763 -0.763 -0.090 0.602 0.784 0.316 0.424 0.316
Gemini-CLI 3.75 4.00 -0.183 -0.356 -0.176 0.000 0.000 -0.995 0.811 -0.193
Qwen-Coder 3.00 2.00 0.632 0.717 0.826 0.705 0.641 0.677 -0.944 0.000
The trend on four games are rather different. In general, the commercial model group consistently demonstrates superior global learning capability, where the advantage is particularly evident in complex strategy games like Chess or Gomoku variant.
Despite most minimal agents fail to learn well on complex games, we still find that Claude-4-Sonnet significantly surpass the rivals on standard games. However, the Claude-4-Sonnet still lacks behind on some cases like Gomoku, indicating that current LLMs agentic ablilities are still limited by the framework, where the commericial agents optimize workflow for their specific models to achieve the best results.
In simple games such as Hold’em, a larger proportion of agents exhibit positive learning scores, whereas in complex games like Chess, the prevalence of negative scores increases markedly. This trend suggests that current agents still face significant limitations in learning complex strategies.
E.3 Behavioral Changes Induced by Learning
For each game, we randomly select 80-100 intermediate states from the agents’ rival history and require the agent or LLM-player to choose the next move for each state. To ensure clarity in our writing, we uniformly refer to these intermediate states as endgame throughout the paper. Please note that endgame here is not limited to the final stages of the game; samples are taken from early, middle, and late stages as well.
We visualize the action consistency among the first two rounds in four games’s endgame in Figure 5.
From the matrix, we observe that, in general, agents tend to learn the strategies of other agents in the first round (lower left part vs. upper left part). Specifically, Doubao-Seed and DeepSeek-Chat simply copy Claude-4-Sonnet’s strategy in Holdem. Additionally, the learning trend is more pronounced in simpler games (Holdem vs. Chess).
For simpler games like Holdem, agent strategies in the second round are more similar to those in the first round (lower right part vs. upper left part), while for more difficult games like Chess, the trend is reversed. This observation is consistent with our findings on $\text{Trend}_{\text{mean}}$ in Table 5.
(a) Holdem.
(b) Gomoku.
(c) Bridge.
(d) Chess.
Figure 5: Action consistency between round 1 and round 2 agents’ code on endgames.
Appendix F Comparison between Agent and LLM-Player
Table 12: Comparison of match outcomes between each agent and its corresponding LLM-Player. Each value indicates the agent’s win rate when competing against the LLM that powers it.
Agent VS LLM Gomoku Hold’em Bridge Chess
Standard Variant Standard Variant Standard Variant Standard Variant
Claude-4-Sonnet 1.00 1.00 0.00 0.00 0.45 1.00 0.88 0.75
Deepseek-Chat 0.50 0.00 0.40 0.30 0.00 0.00 0.00 0.00
Doubao-Seed 0.50 0.25 0.00 0.00 0.00 0.10 0.00 0.00
Gemini-2.5-Pro 0.00 0.25 0.00 0.00 1.00 1.00 0.13 0.50
GPT-5 0.00 0.50 0.00 0.50 1.00 0.00 0.00 0.00
Qwen3-Coder 1.00 0.00 0.50 0.50 1.00 0.00 0.50 0.50
Claude-Code 1.00 1.00 0.00 0.00 0.70 0.85 0.75 0.63
CodeX 0.00 0.50 0.20 0.40 1.00 0.38 1.00 1.00
Gemini-CLI 0.00 0.00 0.10 0.40 0.00 1.00 0.38 0.50
Qwen-Coder 1.00 1.00 0.50 0.00 1.00 0.00 0.56 0.63
We compare the strategies of agents’ code and its corresponding LLM-Player on four games in table 12.
Interestingly, there is no strong correlation between the performance of the agents’ code and that of their underlying models.
In games with strong strategic elements, such as Gomoku and Chess, some agents’ code significantly outperforms their corresponding LLM-Player, indicating that the code implementation is able to better leverage game rules and strategies. For example, the agent developed by claude-4-sonnet achieves a 100% win rate against its LLM-Player in both standard and variant Gomoku, and also demonstrates a high win rate in Chess and Bridge. This suggests that the strategies implemented in the code are superior to the large model’s direct reasoning performance as a player in these games. In contrast, the agents developed by doubao-seed and deepseek-chat struggle to defeat their respective models. However, in Hold’em, agents generally have lower win rates than the LLM-Player, possibly because the LLM-Player performs better in games with more psychological tactics, which are difficult to simulate with code while can be summarized by context learning.
We further visualize the action consistency between agents’ code and LLM-palyers in Figure 6.
We find that, in most games, the actions of LLM-players tend to resemble each other, and the strategies implemented in agents’ code also exhibit high similarity among themselves. However, the code-based strategies and the plain reasoning of the same model often differ substantially. The only exception is Bridge, where we find numerous cases in which both LLM-players and code agents exhibit low consistency with human decisions. Considering that Bridge allows for a certain degree of decision freedom and its bidding rules are not strictly unified, we attribute this phenomenon to the intrinsic characteristics of the game. Similar observations are also reported in other studies (Kita et al., 2024).
These findings further demonstrate that the strategies generated by the agents and those employed by the LLM-players are based on different approaches. This difference merits additional investigation in subsequent studies.
Appendix G Case Study on Code
G.1 Strategy of Agents
In Gomoku, strategies display clear stratification. Gemini and DeepSeek rely on random or near-random moves, while Claude, Doubao, GPT-5, and Qwen3 employ similar pattern-based evaluation with candidate filtering and Minimax search. Differences mainly lie in threat recognition, search control, and opening play: Claude and Doubao handle openings and forced moves more effectively, GPT-5 is steadier under time limits, Qwen3 remains balanced, whereas Gemini and DeepSeek are notably weaker.
The code similarity among agents in Texas Hold’em is relatively high, only DeepSeek employs a fully random strategy, while other models calculate winning probabilities based on the hand. On one hand, this is because the available actions in Texas Hold’em are limited to fold, call/raise, and check. On the other hand, the strategies for Texas Hold’em are relatively straightforward to implement, as both reasoning and code are primarily based on hand strength. As a result, the code can closely simulate the reasoning process.
In the case of Chess, DeepSeek relies on an external library (Stockfish), but fails to configure it correctly, resulting in unsuccessful development. Even after multiple development iterations, DeepSeek continues to use this library without resolving the configuration issues. We also find that Claude, Doubao, Gemini, GPT, and Qwen3 utilize a similar combination of heuristic piece and board evaluation, Minimax search, and alpha-beta pruning, which leads to similar behavior. There are slight differences in how each model evaluates the value of Chess pieces and the actions in endgame scenarios. Notably, Claude incorporates an opening book, which distinguishes it from the others and leads to better performance.
For Bridge, bidding and play strategies also stratify clearly. Qwen3 and Gemini rely on minimal logic, following random choices or only basic rules on High Card Points (HCP). By contrast, GPT-5, Doubao, and Claude incorporate structured evaluation, moving from total point counting (GPT-5) to multi-layered systems with suit quality, competitive actions, and signaling (Doubao and Claude). Despite these differences, all models share reliance on HCP as a core metric. Overall, Claude achieves the most complete integration of evaluation and play, Doubao is comparably advanced, GPT-5 remains simpler, while Gemini and Qwen3 lag behind.
(a) Holdem.
(b) Gomoku.
(c) Bridge.
(d) Chess.
Figure 6: Action consistency between agents’ code and LLM-Players on endgames.
Appendix H Results of ML Track
Table 13: ML ability scores and average rankings of agents.
[…]
Minimal
Claude-4-Sonnet 0.787 0.360 0.600 0.700 2.25
DeepSeek-Chat 0.612 0.000 0.170 0.100 4.25
Doubao-Seed 0.375 0.110 0.290 0.100 4.25
Gemini-2.5-Pro 0.000 0.000 0.140 0.675 5.00
GPT-5 0.625 0.530 0.900 0.700 1.50
Qwen3-Coder 0.600 0.000 0.900 0.725 2.50
Commercial
best ADK 0.750 0.190 0.700 0.656 1.25
Claude-Code 0.578 0.170 0.000 0.406 4.00
CodeX 0.484 0.190 0.400 0.469 3.00
Gemini-CLI 0.187 0.280 0.200 0.438 3.50
Qwen-Coder 0.500 0.170 0.700 0.531 2.50
The detailed results of agents’ performance on machine learning devlopments is shown in table 13.
In ML track, agents autonomously generate data, design code, train models, and deliver ML-based strategies in a GPU-enabled environment. The results of ML ability is provided in Appendix H. Most agents only implement basic models with limited training, resulting in smaller performance gaps and different rankings compared to the strategy track.
Appendix I Results of Multi-lingual Track
We list the detailed results on different languages in  14.
Table 14: Scores of agents on games with different languages, with variance analysis.
[…]
Minimal
Claude-4-Sonnet 1.000 0.250 0.250 0.125 0.360 0.640 0.000 0.069 1.000 0.250 0.250 0.125 0.106
DeepSeek-Chat 1.000 0.250 0.250 0.125 1.000 0.000 0.000 0.222 0.500 0.500 0.500 0.000 0.116
Doubao-Seed 1.000 0.500 0.000 0.167 1.000 0.000 0.000 0.222 0.500 1.000 0.000 0.167 0.185
Gemini-2.5-Pro 0.750 0.750 0.000 0.125 0.010 0.990 0.000 0.216 1.000 0.250 0.250 0.125 0.155
GPT-5 0.500 0.000 1.000 0.167 1.000 0.000 0.000 0.222 1.000 0.500 0.000 0.167 0.185
Qwen3-Coder 1.000 0.250 0.250 0.125 0.610 0.290 0.100 0.044 0.688 0.000 0.812 0.128 0.099
Commercial
Claude-Code 1.000 0.250 0.250 0.125 0.200 0.020 0.780 0.105 1.000 0.250 0.250 0.125 0.118
CodeX 0.000 0.500 1.000 0.167 0.200 0.030 0.770 0.100 1.000 0.000 0.500 0.167 0.145
Gemini-CLI 0.750 0.750 0.000 0.125 1.000 0.000 0.000 0.222 1.000 0.250 0.250 0.125 0.157
Qwen-Coder 1.000 0.250 0.250 0.125 1.000 0.000 0.000 0.222 0.975 0.000 0.525 0.159 0.169
Most agents achieve their highest scores in Python, while several models exhibit significant performance fluctuations in JS and Go. Qwen3-Coder demonstrates the most consistent results across all languages, with the lowest average variance, indicating strong cross-language adaptability. In contrast, models such as GPT-5 and Doubao-Seed show considerable differences between languages, reflecting limited generalization ability. Commercial agents also exhibit score differences across different programming languages. Considering that board game strategies are inherently language-agnostic tasks, the performance gaps observed in the multi-language track of CATArena indicate that current agents are not yet able to effectively abstract strategies into unified algorithms and implement them consistently across languages. Such algorithmic abstraction should be a key direction for the future development of agents.
Appendix J Cost and Code Complexity of Participants
Table 15: Comparison of model cost on four games of 1st round.
Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code CodeX Gemini-CLI Qwen-Coder
Gomoku
Total input tokens 825964 34266 83307 1074278 392378 251590 925166 - 169567 623393
Session tokens 31571 8256 20186 27481 25201 20013 29340 - 21590 26478
Total output tokens 13777 1400 9432 25952 14184 6338 11746 - 2303 8002
Total time (s) 441.916 29.021 206.159 858.027 280.399 135.770 323.200 - 519.100 607.200
Tools used 68 12 20 118 44 32 38 24 8 26
Valid lines of code 574 69 405 65 421 425 350 347 101 335
Avg thinking time (s) 0.0337 0.0026 2.9848 0.0026 1.9681 0.0903 0.0053 0.7026 0.1151 0.0036
Texas Hold’em Poker
Total input tokens 966609.5 44883 97274 168565 345444 326508 296329.5 - 840690.5 1268906.5
Session tokens 45384 9388 14413 14073 26418 16688 29663.5 - 27828 23646.5
Total output tokens 20974 985 10370 40762 24178 4472 8537 - 11833.5 12772
Total time (s) 489.150 25.385 256.145 481.315 473.910 131.630 143.420 - 365.485 474.075
Tools used 38 8 12 15 20 24 12 19 39 48
Valid lines of code 742 21 169 92 346 164 122 318 300 220
Avg thinking time (s) 0.0026 0.0023 0.0030 0.0025 0.1004 0.0025 0.0028 0.0021 0.0024 0.0023
Bridge
Total input tokens 1429350 18027 81206 266506 258633 114679 618991 - 204913 71696
Session tokens 38087 15077 24374 29486 35448 20358 34263 - 46945 34937
Total output tokens 17039 1131 19247 12713 3954 5394 12284 - 13222 20308
Total time (s) 727.550 3.523 333.361 191.775 128.895 334.394 274.200 - 188.700 613.000
Tools used 94 6 14 32 26 16 25 24 5 32
Valid lines of code 521 0 552 280 245 271 464 362 349 549
Avg thinking time (s) 0.0095 0.0000 0.0105 0.0094 0.0095 0.0093 0.0091 0.0091 0.0093 0.0093
Chess
Total input tokens 1612835 49655 86674 346890 484611 267156 709737 - 153825 23489
Session tokens 38729 9953 14799 93292 39962 20781 29818 - 25418 23620
Total output tokens 16584 1257 14411 80997 28695 6624 13017 - 4727 6845
Total time (s) 617.339 26.882 287.976 356.551 444.932 141.417 285.593 - 162.342 142.344
Tools used 110 14 16 40 44 34 32 14 6 18
Valid lines of code 734 0 292 359 437 360 393 283 262 325
Avg thinking time (s) 1.3500 0.0000 0.0030 0.0030 0.2690 0.0020 1.9730 0.0030 0.0030 0.0050
Average across games
Total input tokens 1208689.625 36707.75 87115.25 464059.75 370266.5 239983.25 637555.875 - 342248.875 496871.125
Session tokens 38442.75 10668.5 18443.0 41083.0 31757.25 19460.0 30771.125 - 30445.25 27170.375
Total output tokens 17093.5 1193.25 13365.0 40106.0 17752.75 5707.0 11396.0 - 8021.375 11981.75
Total time (s) 568.989 21.203 270.910 471.917 332.034 185.803 256.603 - 308.907 459.155
Tools used 77.5 10.0 15.5 51.25 33.5 26.5 26.75 20.25 14.375 30.875
Valid lines of code 642.75 22.5 354.5 199.0 362.25 305.0 332.25 327.5 253.0 357.25
Avg thinking time (s) 0.3489 0.0012 0.7503 0.0044 0.5867 0.0260 0.4975 0.1792 0.0324 0.0050
Table 16: Comparison of model cost on four games of 2nd round.
Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code CodeX Gemini-CLI Qwen-Coder
Gomoku
Total input tokens 1924785 170294 625384 589934 1351710 4859091 761468 - 514517 573759
Session tokens 102767 3495 84577 38760 78092 151664 37085 - 53402 27637
Total output tokens 14154 1608 16828 17098 10126 8230 11705 - 5180 7006
Total time (s) 433.112 765.076 436.142 1195.109 387.641 918.349 265.7 - 104.4 273.5
Tools used 58 62 32 58 48 82 30 22 12 23
Valid lines of code 617 69 527 305 467 368 493 302 251 376
Avg thinking time (s) 1.9630 0.0026 2.1291 0.0000 1.5888 0.3499 0.7542 0.2172 0.0080 0.0047
Texas Hold’em Poker
Total input tokens 1497923 56126 225806 2076630 1328520 3671952 1143571 - 1048050 1071708
Session tokens 143575 9819 57890 240057 113649 323278 53929 - 40566 49161
Total output tokens 7309 1151.5 10263.5 46280.5 13269.5 5986.5 7349.5 - 12617 6954
Total time (s) 308.05 45.935 327.225 1171.58 691.74 1057.84 866.715 - 1453.08 163.275
Tools used 16 11 10 17 21 25 28 18 41 27
Valid lines of code 283 21 110 156 362 293 264 319 363 315
Avg thinking time (s) 0.0026 0.0024 0.0024 0.0019 0.0912 0.0026 0.0023 0.0024 0.0025 0.0025
Bridge
Total input tokens 740733 18388 184814 277299 900865 720015 940709 - 275340 105056
Session tokens 28130 15318 4076 131089 119593 62734 38388 - 49946 63756
Total output tokens 14200 1178 18169 7520 4983 10270 15530 - 8161 25125
Total time (s) 397.164 6.097 391.047 162.147 367.204 1863.073 365.9 - 169.0 1337.7
Tools used 48 8 28 26 68 44 33 17 6 38
Valid lines of code 650 0 731 271 245 288 700 367 280 1112
Avg thinking time (s) 0.0101 0.0000 0.0135 0.0099 0.0096 0.0099 0.0105 0.0105 0.0102 0.0106
Chess
Total input tokens 2118018 45051 154263 426659 390211 1014853 4129953 - 308116 31752
Session tokens 101195 3954 28503 26538 35482 36705 72658 - 33919 31877
Total output tokens 13067 412 10063 13435 16798 11018 25438 - 6283 7250
Total time (s) 1352.194 522.132 230.664 931.231 214.792 263.665 906.521 - 420.146 177.453
Tools used 44 26 28 44 34 80 78 14 13 28
Valid lines of code 647 0 305 347 559 499 736 351 299 335
Avg thinking time (s) 0.8240 0.0000 0.0020 0.0000 0.0030 0.0020 0.0030 0.0040 1.8470 0.0050
Average across games
Total input tokens 1570364.75 72464.75 297566.75 842630.5 992826.5 2566477.75 1743925.25 - 536505.75 445568.75
Session tokens 93916.75 8146.5 43761.5 109111.0 86704.0 143595.25 50515.0 - 44458.25 43107.75
Total output tokens 12182.5 1087.38 13830.88 21083.38 11294.13 8876.13 15005.63 - 8060.25 11583.75
Total time (s) 622.63 334.81 346.27 865.02 415.34 1025.73 601.21 - 536.66 487.98
Tools used 41.5 26.75 24.5 36.25 42.75 57.75 42.25 17.75 18.0 29.0
Valid lines of code 549.25 22.5 418.25 269.75 408.25 362.0 548.25 334.75 298.25 534.5
Avg thinking time (s) 0.7000 0.0013 0.5367 0.0029 0.4231 0.0911 0.1925 0.0585 0.4669 0.0057
Table 17: Comparison of model cost on four variant games of 1st round.
Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code CodeX Gemini-CLI Qwen-Coder
Gomoku
Total input tokens 833531 35489 83792 342027 539388 68087 495982 - 256919 23991
Complete tokens 22521 8587 23512 27175 27291 11468 27275 - 39232 24082
Total output tokens 10900 1394 11740 13553 19676 1658 8153 - 3522 7436
Total time (s) 307.319 24.007 255.817 809.100 365.077 73.260 191.800 - 74.300 140.200
Tools used 64 14 20 36 52 14 23 13 7 18
Valid lines of code 859 64 746 414 301 298 254 313 156 431
Avg thinking time (s) 0.0473 0.0026 0.0333 6.7802 0.0904 0.0048 0.0106 0.0051 0.0124 0.7938
Texas Hold’em Poker
Total input tokens 1182600 39782 89270 2582120 191064 385246 160410 - 864568 775880
Complete tokens 39575 9174 14448 32547 16530 16013 24287 - 25654 25308
Total output tokens 13110 962 9897 52457 16407 4361 5801 - 10699 7568
Total time (s) 398.130 20.195 217.600 1337.345 558.055 124.640 84.920 - 319.745 270.315
Tools used 42 7 11 86 17 30 7 19 42 34
Valid lines of code 469 21 204 98 302 144 167 256 232 243
Avg thinking time (s) 0.0026 0.0022 0.0011 0.0486 0.0019 0.0011 0.0025 0.0801 0.0028 0.0021
Bridge
Total input tokens 475377 16805 122427 232284 110768 653245 1908005 - 173628 26896
Complete tokens 37640 13764 36846 22593 4433 4438 53465 - 33184 27023
Total output tokens 16704 1137 7583 10020 2378 22823 18489 - 7412 5191
Total time (s) 298.800 7.721 158.254 483.948 304.010 7111.237 462.300 - 229.200 129.600
Tools used 34 6 16 32 40 308 55 16 5 22
Valid lines of code 627 0 291 213 133 165 469 368 212 224
Avg thinking time (s) 0.0071 0.0000 0.0068 0.0068 0.0096 0.0067 0.0089 0.0089 0.0060 0.0092
Chess
Total input tokens 1000994 38977 84149 432950 229279 253674 979611 - 184178 115022
Complete tokens 38485 7761 17321 43002 20520 17592 34597 - 26984 19923
Total output tokens 16355 1125 10518 32332 16499 5948 13010 - 6014 8032
Total time (s) 473.286 23.186 241.221 729.121 283.023 144.435 357.587 - 139.114 368.952
Tools used 74 14 14 40 32 38 41 12 8 19
Valid lines of code 719 41 399 398 475 343 410 427 152 392
Avg thinking time (s) 0.7520 0.0000 0.0000 0.2560 1.1190 0.0020 1.9660 0.0040 0.0040 0.0050
Average across games
Total input tokens 873125.5 32763.25 94909.5 897345.25 267624.75 340063.0 886002.0 - 369823.25 235447.25
Complete tokens 34555.25 9821.5 23031.75 31329.25 17193.5 12377.75 34906.0 - 31263.5 24084.0
Total output tokens 14267.25 1154.5 9934.5 27090.5 13740.0 8697.5 11363.25 - 6911.75 7056.75
Total time (s) 369.3838 18.7773 218.2230 839.8785 377.5413 1863.3930 274.1518 - 190.5898 227.2668
Tools used 53.5 10.25 15.25 48.5 35.25 97.5 31.5 15.0 15.5 23.25
Valid lines of code 668.5 31.5 410.0 280.75 302.75 237.5 325.0 341.0 188.0 322.5
Avg thinking time (s) 0.2022 0.0012 0.0103 1.7729 0.3052 0.0036 0.4970 0.0245 0.0063 0.2025
Table 18: Comparison of model cost on four variant games of 2nd round.
Metric Claude-4-Sonnet Deepseek-Chat Doubao-Seed Gemini-2.5-Pro GPT-5 Qwen3-Coder Claude-Code CodeX Gemini-CLI Qwen-Coder
Gomoku
Total input tokens 1298765 111940 680676 545504 606704 1542046 805252 - 227722 873934
Complete tokens 132219 11100 121218 24003 82812 4309 38629 - 41093 32599
Total output tokens 7118 1902 26565 18241 4189 22014 11529 - 5554 8531
Total time (s) 261.323 76.973 538.087 325.892 179.801 25787.922 294.400 - 80.200 836.300
Tools used 46 30 28 70 38 538 30 23 6 30
Valid lines of code 264 64 535 111 276 43 328 302 210 384
Avg thinking time (s) 0.0212 0.0030 0.0855 0.0000 0.0904 0.0000 0.0884 0.0162 0.0070 0.0071
Texas Hold’em Poker
Total input tokens 1627500 50339 656196 675580 1452101 1822342 448924 - 1725678 2472353
Complete tokens 143514 11897 115933 103332 111407 118321 48631 - 48378 67453
Total output tokens 8180 1130 11609 18771 12897 6227 5880 - 16033 10296
Total time (s) 367.720 27.960 359.490 515.000 408.505 294.910 612.525 - 670.535 283.440
Tools used 18 10 16 20 22 27 16 23 53 48
Valid lines of code 289 22 272 105 324 295 167 256 232 243
Avg thinking time (s) 0.0025 0.0019 0.0019 0.0031 0.0019 0.0026 0.0025 0.0801 0.0028 0.0021
Bridge
Total input tokens 706814 17186 32819 775688 433901 866917 1110504 - 189659 48922
Complete tokens 41164 14021 14545 80161 $41$ 4944 45001 - 46073 49034
Total output tokens 15867 1188 13212 39919 1485 13928 17233 - 7966 7252
Total time (s) 359.864 6.074 189.645 279.231 1106.275 3238.660 411.500 - 77.700 220.000
Tools used 52 8 12 46 112 158 38 21 4 25
Valid lines of code 603 0 233 396 166 218 747 295 441 531
Avg thinking time (s) 0.0069 0.0000 0.0112 0.0067 0.0065 0.0101 0.0093 0.0091 0.0091 0.0092
Chess
Total input tokens 1475420 34542 172443 400275 473325 1682584 2213785 - 247641 109313
Complete tokens 56660 8283 31582 97638 32082 44129 43227 - 23389 8140
Total output tokens 22989 1473 16610 35371 16469 14241 24090 - 4834 9129
Total time (s) 696.717 32.412 366.723 1093.706 350.225 511.943 605.236 - 237.026 427.591
Tools used 82 14 30 62 44 116 67 16 13 29
Valid lines of code 1146 72 561 345 506 577 407 434 187 454
Avg thinking time (s) 1.6260 0.0000 0.0000 0.0440 0.7560 0.1780 0.0030 0.0040 1.7130 0.0050
Average across games
Total input tokens 1277124.75 53501.75 385533.5 599261.75 741507.75 1478472.25 1144616.25 - 597675.0 876130.5
Complete tokens 93389.25 11325.25 70819.5 76283.5 57610.5 42925.75 43872.0 - 39733.25 39306.5
Total output tokens 13538.5 1423.25 16999.0 28075.5 8760.0 14102.5 14683.0 - 8596.75 8802.0
Total time (s) 421.4060 35.8547 363.4863 553.4572 511.2015 7458.3587 480.9153 - 266.3653 441.8328
Tools used 49.5 15.5 21.5 49.5 54.0 209.75 37.75 20.75 19.0 33.0
Valid lines of code 575.5 39.5 400.25 239.25 318.0 283.25 412.25 321.75 267.5 403.0
Avg thinking time (s) 0.4141 0.0012 0.0247 0.0134 0.2137 0.0477 0.0258 0.0273 0.4330 0.0059
We list the agents’ cost and code statistics in table 15 for first round development of standard games,  16 for second round development of standard games, 17 for first round development of vairant games and  18 for second round development of variant games.
We can see that game development token costs show minimal variation, while differences is significant due to model changes. Claude (both minimal and code-based agents) consumes significantly more input tokens than competitors, exceeding the average by over 2 times, while Gemini generates notably more output tokens compared to other models. GPT-5 offers the best trade-off between cost. Among all agents, second-round game development require more input tokens, while output token growth remains marginal. In addition, commercial agents consistently use fewer tokens than their minimal-agent counterparts.
In terms of code complexity, agents driven by Claude-4 model consistently surpasses other agents in both the number of effective lines of code developed and the time spent considering development strategies. We observe that its development strategies are more sophisticated. Additionally, the complexity of its code increases with each iteration, which indirectly demonstrates the model’s exceptional learning capabilities.
Appendix K Full Prompts
The agent is instructed to develop a competitive game AI based on the provided game environment. The AI must be deployed as an HTTP service with a single-port startup script, follow the official development instructions, and be named with the model prefix. And the agent is encouraged to iteratively improve its strategy based on the tournament report and the previous codes. Full prompt of details are in Table 19 for main leaderboard,  LABEL:tab:ml_game_ai_prompt for ML track and  21 for multilingual track.
Table 19: Game AI Development Prompt
Game Development
There is a game project under game env path. You need to read its source code and develop a game AI. Your AI will compete against other AIs in a tournament, so please make your AI as strategic and competitive as possible.
The final AI should be provided as an HTTP service. You can refer to the guides in game env path / README.md and game env path / develop instruction.md for development instructions. The content in game env path / develop instruction.md is very important, please read it carefully!
Please develop your AI service directly under dir path.
Script Requirements
Please implement a script to start your AI service, with the name start_ai.sh in dir path. The script must accept exactly one argument, which is the port number game port to run the HTTP service. You should be able to start the AI service on a specified port by running:
bash start_ai.sh <port>
Your AI service should listen on the given port, and you can check its health status by running: curl -s http://localhost:<port>/health
Note: The script should not accept any other arguments except for the port number. Make sure your AI service uses this port for HTTP requests.
Other Requirements:
Use your model name as a prefix in the name of your AI service, i.e., model name_AI. Develop directly in dir path without repeatedly asking for the next step. Report to me only after you have completed the development.
Access the main server
You can play game of game env pathin at game server. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
Final Remind
You should write start_ai.sh in dir path and implement the AI service in dir path. DO NOT MODIFY THE CODE IN game env path
Condition (if round_num>1round\_num>1):
Tournament report of last round is in last round log dir. The historical records are quite large, please use tools start_interactive_shell and run_interactive_shell to analyze the data efficiently. The code corresponding to the log is stored in last round code dir. Please learn from it and improve your strategy.
Table 20: Machine Learning Game AI with MANDATORY Self-Play Training Prompt
Machine Learning Game AI with MANDATORY Self-Play Training
Develop a competitive game AI for game env path using REAL machine learning with actual training.
CRITICAL: NO PSEUDO-ML ALLOWED
MANDATORY: Implement real training with actual parameter updates.
Forbidden: Random weights, unused optimizers, no training loops
Required: Self-play training, loss.backward(), optimizer.step(), saved trained model
Training Requirements
1. Self-Play System: Generate training data by playing against itself
2. Training Loop: Real parameter updates with backpropagation
3. Model Saving: Save trained model weights (e.g., trained_model.pth)
4. Training Endpoint: /train HTTP endpoint to trigger training
Technical Implementation
The final AI should be provided as an HTTP service. You can refer to the guides in game env path/README.md and game env path/develop_instruction.md for development instructions. The content in game env path/develop_instruction.md is very important, please read it carefully!
Please develop your AI service directly under dir path.
Script Requirements
Implement a script to start your AI service, with the name start_ai.sh in dir path. The script must accept exactly one argument, the port number to run the HTTP service. Start the AI via:
bash start_ai.sh <port>
Check service health via: curl -s http://localhost:<port>/health
Additional Training Script: Also create train_ai.sh for self-play training:
bash train_ai.sh <num_episodes>
Note: The script should not accept any other arguments except for the port number. Ensure the AI service uses this port for HTTP requests.
Environment Requirements
MANDATORY: You MUST develop under the ml_env_py311 conda environment. PyTorch, TensorFlow, NumPy, pandas, and scikit-learn are already installed.
- DO NOT include PyTorch or TensorFlow in requirements.txt
- Create a requirements.txt file only for additional dependencies you need beyond the pre-installed packages
- All development and execution must be done in the ml_env_py311 environment
Available Game Port
We offer an interactive game port: http://localhost:game port to facilitate your testing and training. You can interact with the service and get feedback from the port.
Validation Requirements
Submission will be rejected if:
- The model uses only random weights without training
- No actual training loop with parameter updates
- No self-play data generation system
- Cannot demonstrate learning/improvement over time
- Training endpoints return fake responses without real training
Other Requirements
Use your model name as a prefix in the name of your AI service, i.e., model name_AI.
Develop directly in dir path without repeatedly asking for the next step. Report to me only after you have completed the development.
Final Remind
You should write start_ai.sh and train_ai.sh in dir path and implement the REAL ML-based AI service with actual training capabilities in dir path. DO NOT MODIFY THE CODE IN game env path. Please make sure to implement a genuine machine learning training workflow, including writing the code, performing training, tuning hyperparameters, and testing the model. After training, save the trained model parameters, and ensure that the AI service can be started with the specified script to perform inference.
Mandatory: Submission must include a trained model file (e.g., trained_model.pth, model_weights.pkl) proving real training.
Zero Tolerance for Pseudo-ML: Any submission without real training will be rejected.
Table 21: Multi-language Game AI Development Prompt
Game Development
There is a game project under game env path. You need to read its source code and develop a game AI. Your AI will compete against other AIs in a tournament, so please make your AI as strategic and competitive as possible.
The final AI should be provided as an HTTP service. You can refer to the guides in game env path / README.md and game env path / develop instruction.md for development instructions. The content in game env path / develop instruction.md is very important, please read it carefully!
Please develop your AI service directly under dir path.
Script Requirements
Please implement a script to start your AI service, with the name start_ai.sh in dir path. The script must accept exactly one argument, which is the port number game port to run the HTTP service. You should be able to start the AI service on a specified port by running:
bash start_ai.sh <port>
Your AI service should listen on the given port, and you can check its health status by running: curl -s http://localhost:<port>/health
Note: The script should not accept any other arguments except for the port number. Make sure your AI service uses this port for HTTP requests.
Other Requirements:
Use your model name as a prefix in the name of your AI service, i.e., model name_AI. Develop directly in dir path without repeatedly asking for the next step. Report to me only after you have completed the development.
Access the main server
You can play game of game env pathin at game server. You can play the games with your own AI or any other AI to improve your strategy. You can use bash tools to improve yourself.
Final Remind
You should write start_ai.sh in dir path and implement the AI service in dir path. DO NOT MODIFY THE CODE IN game env path
Condition (if language = JS)
JavaScript is the language you should use to develop your AI service. The version of Node.js is node version, the path of Node.js is node path, and it is already set in the PATH environment variable. You can use node to run the program.
Condition (if language = Go)
Go is the language you should use to develop your AI service. The version of Go is go version, the path of Go is go path, and it is already set in the PATH environment variable. You can use go to build the program.